# references.bib

### v2

@@ 6. When can a decision-maker be excused for unlawful discrimination?
6.1. Direct Discrimination has no defence
A unique feature of UK direct discrimination is that: “In contrast to the law in many countries, where English law forbids direct discrimination it provides no defence of justification” (EandJFS). In the UK Supreme Court (UKSC), Lord Phillips explained that while “it is possible to envisage circumstances where giving preference to a minority racial [or other protected] group will be justified” nevertheless “a policy which directly favours one racial [or other protected] group will be held to constitute racial discrimination against all who are not members of that group” (EandJFS). Lady Hale confirmed that “however justifiable it might have been, however benign the motives of the people involved, the law admits of no defence” (EandJFS). Given direct algorithmic discrimination is more likely to arise in the UK, the absence of a defence is significant. It is important to correct misconceptions that direct discrimination are either unlikely to arise or may be excused, such as in the US where a legitimate, non-discriminatory justification can excuse disparate treatment.
6.2. Indirect discrimination may be justified
On the other hand, indirect discrimination can be excused if the PCP is objectively justified as a proportionate means of achieving a legitimate aim (EqualityAct, s 19(2)(d)). Decision-makers must consider the legitimacy of using an SML model by explicitly defining its purpose and the outcome variable $y$. In algorithm design, social implications should be considered (mulligan2019; hu2020; kasy2021), as well as alignment with legal expectations.
@@ 6.2.1. Defining legitimate aims $y$
Identifying a legitimate aim is closely connected to the choice of $y$, the unknown entity used for decision-making. The legitimacy of the aim depends on the decision-makers’ raison d’être (khaitan2015; justice1986; court2006). In Homer, the UKSC established a legitimate aim must “correspond to a real need and the means used must be appropriate with a view to achieving the objective and be necessary to that end” (court2012). For example, in lending, it is a legitimate aim to protect the repayment of their loans or at least secure their loans. In fact, “the mortgage market could not survive without that aim being realised” (bristol2015). If the choice of $y$ is legitimate based on context and the benefit outweighs any potential harm, there is a lower risk of unlawful discrimination (justice1986).
For a legitimate $y$ to be an exception to indirect discrimination, the PCP must be a proportionate means of achieving the legitimate $y$ (EqualityAct). To be proportionate, it must be an appropriate means of achieving the legitimate aim and (reasonably) necessary to do so (court2012). Such analysis will turn on the facts of each case by evaluating whether the design choices were “appropriate with a view to achieving the objective and be necessary” and weighing the need against the seriousness of detriment to the disadvantaged group (court2006). Proportionality assessments balance the discriminatory effect of a PCP with the reasonable needs of a business (lane2018), and consider whether non-discriminatory alternatives were available (court2012). While important work on identifying less discriminatory models is emerging (finreglab_machine_2023; gillis2024; black2024), such efforts should be viewed within the broader context. A less discriminatory algorithm should not substitute pursuing less discriminatory alternatives, which may involve dispensing with algorithmic decision-making altogether. Measures to improve accuracy, maximise benefits over costs, minimise estimation error, or condition for protected attributes may all be relevant considerations for whether the modeller’s choices were proportionate means of achieving a legitimate $y$.
@@
To illustrate with an example, if a bank’s training data is outdated or sourced from a different country, it may not accurately represent the current population relevant to the model. This discrepancy can lead to biased estimates, particularly if the data reflects historical prejudices. For instance, the model might unjustly associate certain demographics with higher default risk, not because of true differences in $y$ but biased data in $\tilde{y}$ (as warned in dfs2021).
@@ 6.2.2. Defining legitimate and non-legitimate variables $x$
For example, in a lending example, hair length strongly correlates to gender in many cultural contexts but is unlikely to contribute to a consumer’s true default risk. Boyarskaya et al. explain the absence of a “causal story” between hair length and loan repayment because hair length would not be part of a true model for the risk of default (boyarskaya2022). Therefore, hair length is an example of $x_{n}$ in a lending context.
For comparison, the legitimacy of zip codes illustrates the nuanced nature of legitimate features. While a zip code may correlate with race in some contexts, it might be a legitimate variable in other situations. For example, in an application for home insurance covering flood risk, zip codes are invaluable proxies for granular information such as geographical features, land topography, and historical flooding. Therefore, in the best model for property flood insurance decisions, zip code will improve the predictive performance as a legitimate proxy for the true geographical features within the true DGP. However, in a university application, zip code should not be predictive or causal to a prospective student’s merit for acceptance. In such cases, zip code likely acts as a proxy for race or the unprotected characteristic of socio-economic status and would be $x_{n}$. So, in some circumstances, the zip code would be legitimate $x_{l}$, but in others, it may be $x_{n}$. It will also be relevant to consider whether a less discriminatory feature is available, i.e., one with less correlation to a protected attribute that is equally predictive.
As explored in Section 7, in lending predictions information about income, employment, and debts are likely to be legitimate features $x_{l}$. Credit scores, or related features, would have a material impact on the true model for default, and then would be a legitimate feature $x_{l}$ (black1978; hurley2017). Given that nearly all features may contain some information on protected attributes, even legitimate factors (corbett2023), this approach explains the need to assess the strength of this dependence and whether the feature contributes significantly to the model’s prediction and can be argued to be part of a true DGP.
@@ 7. Case Study
To demonstrate a real-world approach to the legal reasoning through our formalisation, we now discuss our framework by reference to the first case regarding automated decision-making and discrimination, decided by the National Non-Discrimination and Equality Tribunal of Finland (Tribunal(finland). Finnish anti-discrimination law bears many similarities to UK and EU laws. In Appendix A, we include the relevant provisions of the Finnish Non-Discrimination Act to demonstrate the similarities to the UK Equality Act.
In this case, a person (A), was denied credit for online purchases based on a credit rating system employed by a bank. The Tribunal found that the bank’s statistical scoring model resulted in discrimination based on multiple protected characteristics and was not justified by an acceptable objective achieved by proportionate measures. Consequently, the Tribunal prohibited the bank from continuing this practice and imposed a conditional fine to enforce compliance.
@@ 7.1. Decision-making Model and Data
The decision-making system in question is for online store financing credit. The credit applied for by the consumer in each situation is also always bound to the purchase and its value, which means that it is more difficult, or even impossible, to undertake detailed requests for information and background checks. The company made credit decisions based on data from the credit company’s internal records, credit file information, and the score from the company’s internal scoring system. The bank’s scoring system assessed creditworthiness. The scoring system used population statistics and personal attributes to calculate the percentage of people in certain groups with bad credit history and awarded points proportionate to how common bad credit records were in the group in question. The variables used included race, first language, age, and place of residence. The company did not require or investigate the applicant’s income or financial situation.
@@ 7.2. True Data Generating Process and Estimation Error
The bank’s scoring model was based on multiple variables where the majority was variables using protected attributes, including gender, language, age and place of residence, meaning the model is more or less $\hat{p}(y|x_{p})$.
This model did not attempt to model the true underlying DGP and instead relied on data that was available. Therefore, the model lacks information that could have been used to make better predictions, i.e., legitimate features $x_{l}$. By solely using the data available, rather than identifying data that would be best to reduce conditional estimation error, the modellers built an automated decision-making system that unlawfully discriminated. We now evaluate how the Tribunal came to those conclusions about the legitimacy of $y$ and $x$ for such a model.
@@ 7.3. Legitimate $y$
However, the Tribunal clarified that the creditworthiness assessment “means expressly the assessment of an individual’s credit behaviour, credit history, income level and assets, and not the extension of the impact of models formed on the basis of probability assessments created with statistical methods using the behaviour and characteristics of others, to the individual applying for the credit in the credit decision in such a way that assessment is solely based on such models.” To be appropriate and necessary to achieve that aim, the model must use legitimate features $x_{l}$.
@@ 7.4. Legitimate, Non-Legitimate, and Protected Variables $x$
Four protected characteristics $x_{p}$ were used as variables in this model. The Tribunal considered whether the use of these protected attributes was legitimate.
Age was a non-legitimate variable in this context, but the Tribunal acknowledged that age may be a legitimate variable if it had been used in the assessment of creditworthiness of young persons. Age did not contribute to the accuracy of the model’s prediction in a way that could be argued as part of the true DGP.
First language was a non-legitimate variable in the credit assessment because it “will result, de facto, in the segregation on ethnic lines, the justification of which does not include compelling arguments that could be deemed acceptable from the point of view of the system of fundamental rights.” Evidence showed the model ranked Finnish-speaking residents lower compared to Swedish-speaking residents. Further, ethnic minorities with an official first language other than Finnish or Swedish were put in a more unfavourable position.
Place of residence was a non-legitimate variable because the bank had not proven anything to support an “assumption made on the basis of the general data of residences in a certain area would prove anything of the loan repayment capacity of an individual resident in that area.”
Gender could not be a legitimate variable because it is prohibited from being used as an actuarial factor in financial services under EU law (TestAchats). The express legal prohibition, in this context, means the Tribunal considered that gender should not be part of the true DGP. Although it was shown that women received a higher score than men and that if Person A had been a woman, he would have been granted the credit.
@@ 7.5. Conditional Estimation Parity
Building upon the legitimate variables identified above, we can now examine the concept of conditional estimation parity, which pertains to the disparity in estimation error between groups distinguished by a protected attribute, conditional on legitimate features. Reducing the estimation error articulated in Eq. 4 would mitigate the risk of conditional estimation disparity. However, assessing conditional estimation parity is affected by the inherent challenges in evaluating estimation error. In this case, the Ombudsman brought evidence of the effects of the protected characteristics $x_{p}$ on the true prediction. For example, evidence showed that if conditioned for other variables the model still ranked Finnish-speaking and Swedish-speaking residents differently, meaning conditional estimation parity was not met. The relevant legal evaluation is then whether it is a valid true difference or, as in this case, it is based on a protected characteristic $x_{p}$. Judicial legal reasoning is complex and requires engaging with this type of reasoning through statistical or theoretical means. Unlike other work that has proposed quantitive thresholds or metrics for legality, we focus on formalising modelling choices within UK anti-discrimination law. This approach allows us to outline a flexible, decision-theoretic approach where modeller and decision-maker choices are framed as optimisation under legal constraints, aligning more closely with judicial reasoning while still offering a structured, legally informed framework.
@@ 8. Conclusion
This paper contributes to the understanding of unlawful discrimination in SML under UK law and related jurisdictions. Unlike the statistical focus in much of the algorithmic fairness literature, we present a legally grounded, decision-theoretic framework centred on the true DGP and its connection to legal causation. By emphasising the legitimacy of prediction targets and features, we demonstrate how legally informed concepts, like conditional estimation parity, offer a more cohesive means of assessing when and how models discriminate. By situating these considerations within a UK context that shares key principles with many other jurisdictions, our approach enables more robust global approach to legal doctrine. Crucially, our work corrects several mischaracterisations around discrimination, including those stemming from (a) literature that misrepresents US law, fostering assumptions in the field that are not legally-informed, and (b) literature that accurately reflects US law but diverges materially from UK law.

### v3

@@ 6. When can a decision-maker be excused for unlawful discrimination?
6.1. No Defence for Direct Discrimination
A unique feature of UK direct discrimination is that: “In contrast to the law in many countries, where English law forbids direct discrimination it provides no defence of justification” (United Kingdom Supreme Court, 2009). In the UKSC, Lord Phillips explained that while “it is possible to envisage circumstances where giving preference to a minority racial [or other protected] group will be justified” nevertheless “a policy which directly favours one racial [or other protected] group will be held to constitute racial discrimination against all who are not members of that group” (United Kingdom Supreme Court, 2009). Lady Hale confirmed that “however justifiable it might have been, however benign the motives of the people involved, the law admits of no defence” (United Kingdom Supreme Court, 2009). Given direct algorithmic discrimination is more likely to arise in the UK, the absence of a defence is significant. It is important to correct misconceptions that direct discrimination are either unlikely to arise or may be excused, such as in the US where a legitimate, non-discriminatory justification can excuse disparate treatment.
6.2. Potential Justification of Indirect Discrimination
On the other hand, indirect discrimination can be excused if the PCP is objectively justified as a proportionate means of achieving a legitimate aim (United Kingdom Parliament, 2010, s 19(2)(d)). Decision-makers must consider the legitimacy of using an SML model by explicitly defining its purpose and the outcome variable $y$. In algorithm design, social implications should be considered (Mulligan et al., 2019; Hu and Kohler-Hausmann, 2020; Kasy and Abebe, 2021), as well as alignment with legal expectations.
@@ 6.2.1. Defining legitimate aims $y$
Identifying a legitimate aim is closely connected to the choice of $y$, the unknown entity used for decision-making. The legitimacy of the aim depends on the decision-makers’ raison d’être (Khaitan, 2015; European Court of Justice, 1986; England and Wales Court of Appeal, 2006). In Homer, the UKSC established a legitimate aim must “correspond to a real need and the means used must be appropriate with a view to achieving the objective and be necessary to that end” (United Kingdom Supreme Court, 2012). For example, in lending, it is a legitimate aim to protect the repayment of their loans or at least secure their loans. In fact, “the mortgage market could not survive without that aim being realised” (United Kingdom County Court (2015), Bristol). If the choice of $y$ is legitimate based on context and the benefit outweighs any potential harm, there is a lower risk of unlawful discrimination (European Court of Justice, 1986).
For a legitimate $y$ to be an exception to indirect discrimination, the PCP must be a proportionate means of achieving the legitimate $y$ (United Kingdom Parliament, 2010). To be proportionate, it must be an appropriate means of achieving the legitimate aim and (reasonably) necessary to do so (United Kingdom Supreme Court, 2012). Such analysis will turn on the facts of each case by evaluating whether the design choices were “appropriate with a view to achieving the objective and be necessary” and weighing the need against the seriousness of detriment to the disadvantaged group (England and Wales Court of Appeal, 2006). Proportionality assessments balance the discriminatory effect of a PCP with the reasonable needs of a business (Lane and Ingleby, 2018), and consider whether non-discriminatory alternatives were available (United Kingdom Supreme Court, 2012). While important work on identifying less discriminatory models is emerging (Blattner and Spiess, 2023; Gillis et al., 2024; Black et al., 2024), such efforts should be viewed within the broader context. A less discriminatory algorithm should not substitute pursuing less discriminatory alternatives, which may involve dispensing with algorithmic decision-making altogether. Measures to improve accuracy, maximise benefits over costs, minimise estimation error, or condition for protected attributes may all be relevant considerations for whether the modeller’s choices were proportionate means of achieving a legitimate $y$.
@@
The approximation of the true target variable $\tilde{y}$ can introduce target-construct mismatch, meaning the proxy target variable does not fully align with the underlying phenomenon it aims to measure (Jacobs and Wallach, 2021; Wang et al., 2024). To illustrate with an example, if a bank’s training data is outdated or sourced from a different country, it may not accurately represent the current population relevant to the model. This discrepancy can lead to biased estimates, particularly if the data reflects historical prejudices. For instance, the model might unjustly associate certain demographics with higher default risk, not because of true differences in $y$ but biased data in $\tilde{y}$ (as warned in DFS, 2021).
@@ 6.2.2. Defining legitimate and non-legitimate variables $x$
Traditional guidance often advocates incorporating all available data to maximise predictive accuracy, typically without explicit consideration of causal links between model inputs and the predicted outcomes (Berk, 2019; Manski, 2022; Manski et al., 2023). However, in the presence of label bias, measurement error, or other issues arising from flawed data-generating processes, these approaches risk introducing or amplifying algorithmic bias (Obermeyer et al., 2019; Jacobs and Wallach, 2021; Zanger-Tishler et al., 2024; Mikhaeil et al., 2024). In a medical context, Obermeyer et al. (2019) highlight the importance of modifying the data labels provided to algorithms, emphasising that this process demands deep domain-specific knowledge, iterative experimentation, and careful selection of labels that genuinely represent the true outcomes of interest. For legally compliant models, it is crucial to move beyond naïve performance maximisation toward more nuanced, context-sensitive approaches. We argue that deliberate feature selection should explicitly reflect the causal relationships between model inputs, the target variable, and the underlying true DGP.
For example, in lending, hair length strongly correlates to gender in many cultural contexts but is unlikely to contribute to a consumer’s true default risk. Boyarskaya et al. (2022) explain the absence of a “causal story” between hair length and loan repayment because hair length would not be part of a true model for the risk of default. Therefore, hair length is an example of $x_{n}$ in a lending context.
For comparison, the legitimacy of zip codes illustrates the nuanced nature of legitimate features. While a zip code may correlate with race in some contexts, it might be a legitimate variable in other situations. For example, in an application for home insurance covering flood risk, zip codes are invaluable proxies for granular information such as geographical features, land topography, and historical flooding. Therefore, in the best model for property flood insurance decisions, zip code will improve the predictive performance as a legitimate proxy for the true geographical features within the true DGP. However, in a university application, zip code should not be predictive or causal to a prospective student’s merit for acceptance. In such cases, zip code likely acts as a proxy for race or the unprotected characteristic of socio-economic status and would be $x_{n}$. So, in some circumstances, the zip code would be legitimate $x_{l}$, but in others, it may not be legitimate $x_{n}$. It will also be relevant to consider whether a less discriminatory feature is available, i.e., one with less correlation to a protected attribute that is equally predictive.
As explored in Section 7, in lending predictions information about income, employment, and debts are likely to be legitimate features $x_{l}$. Credit scores, or related features, would have a material impact on the true model for default, and then would be a legitimate feature $x_{l}$ (Black et al., 1978; Hurley and Adebayo, 2017). Given that nearly all features may contain some information on protected attributes, even legitimate factors (Corbett-Davies et al., 2023), this approach explains the need to assess the strength of this dependence and whether the feature contributes significantly to the model’s prediction and can be argued to be part of a true DGP.
@@ 7. Case Study
To demonstrate a real-world approach to the legal reasoning through our formalisation, we now discuss our framework by reference to the first case regarding automated decision-making and discrimination, decided by the National Non-Discrimination and Equality Tribunal of Finland (Tribunal(Finland National Non-Discrimination and Equality Tribunal, 2018). At the time of writing, however, there are no reported UK court decisions that scrutinise algorithmic discrimination in an equivalent setting, and only a handful of such judgments exist worldwide (even fewer outside the United States). Given Finnish anti-discrimination law bears many similarities to UK and EU laws, we use this comparative case analysis to concretely illustrate our framework rather than hypothesise. In Appendix A, we include the relevant provisions of the Finnish Non-Discrimination Act to demonstrate the similarities to the UK Equality Act.
In this case, Person A, was denied credit for online purchases based on a credit rating system employed by a bank. The Tribunal found that the bank’s statistical scoring model resulted in discrimination based on multiple protected characteristics and was not justified by an acceptable objective achieved by proportionate measures. Consequently, the Tribunal prohibited the bank from continuing this practice and imposed a conditional fine to enforce compliance.
@@ 7.1. Decision-Making Model and Data
The decision-making system in question is for online store financing credit. The credit applied for by the consumer in each situation is also always bound to the purchase and its value, which means that it is more difficult, or even impossible, to undertake detailed requests for information and background checks. The credit decisions were based on data from the credit company’s internal records, credit file information, and the score from the company’s internal scoring system. The bank’s scoring system assessed creditworthiness. The scoring system used population statistics and personal attributes to calculate the percentage of people in certain groups with bad credit history and awarded points proportionate to how common bad credit records were in the group in question. The variables used included race, first language, age, and place of residence. The scoring system did not require or investigate the applicant’s income or financial situation.
@@ 7.2. True Data Generating Process and Estimation Error
The bank’s scoring model was based on multiple variables where the majority were protected attributes, including gender, language, age and place of residence, meaning the model is more or less $\hat{p}(y|x_{p})$.
This model did not attempt to model the true underlying DGP and instead relied on data that was available. It is reasonable to expect that the bank was aware of other legitimate factors that could explain the credit score. Therefore, the model lacks information that could have been used to make better predictions, i.e., legitimate features $x_{l}$. By solely using the data available, rather than identifying data that would be best to reduce conditional estimation error, the modellers built an automated decision-making system that unlawfully discriminated. We now evaluate how the Tribunal came to those conclusions about the legitimacy of $y$ and $x$ for such a model.
@@ 7.3. Legitimate $y$
However, the Tribunal clarified that the creditworthiness assessment “means expressly the assessment of an individual’s credit behaviour, credit history, income level and assets, and not the extension of the impact of models formed on the basis of probability assessments created with statistical methods using the behaviour and characteristics of others, to the individual applying for the credit in the credit decision in such a way that assessment is solely based on such models.” To be appropriate and necessary to achieve that aim, therefore, the model must use legitimate features $x_{l}$.
@@ 7.4. Legitimate, Non-Legitimate, and Protected Variables $x$
The Tribunal evaluated each input variable against two cumulative tests: (i) its factual relevance to an individual’s likelihood of repayment and (ii) its permissibility under fundamental-rights and sector-specific law. Variables that clear both hurdles can be labelled legitimate ($x_{l}$); those that fail the predictive-relevance test are non-legitimate ($x_{n}$); and those barred by anti-discrimination law, even if potentially predictive, form the set of protected variables ($x_{p}$).
@@
Four protected characteristics $x_{p}$ were used as variables in this model. The Tribunal considered whether the use of these protected attributes was legitimate. Unlike the economic variables listed above, the following attributes either lacked a demonstrable causal link to repayment or were expressly prohibited by law.
Age was a non-legitimate variable in this context, but the Tribunal acknowledged that age may be a legitimate variable if it had been used in the assessment of creditworthiness of young persons with limited credit history. Age did not contribute to model accuracy in a way that could be argued as part of the true DGP.
First language was a non-legitimate variable in the credit assessment because it “will result, de facto, in the segregation on ethnic lines, the justification of which does not include compelling arguments that could be deemed acceptable from the point of view of the system of fundamental rights.” Evidence showed the model ranked Finnish-speaking residents lower than Swedish-speaking residents. Further, ethnic minorities with another official first language were put in a more unfavourable position.
Place of residence was a non-legitimate variable because the bank had not provided any empirical evidence that an “assumption made on the basis of the general data of residences in a certain area would prove anything of the loan repayment capacity of an individual resident in that area.” Absent such evidence, postcode information remains no more than a group-level stereotype and therefore fails both Tribunal tests.
Gender could not be a legitimate variable because it is prohibited from being used as an actuarial factor in financial services under EU law (European Court of Justice, 2011), also implemented in UK law (United Kingdom Parliament, 2012b). The express legal prohibition, in this context, means the Tribunal considered that gender should not be part of the true DGP. Although it was shown that women received a higher score than men and that if Person A had been a woman, he would have been granted the credit. Therefore, the Tribunal’s approach demonstrates that statistical disparities alone cannot override normative and legal prohibitions.
@@ 7.5. Conditional Estimation Parity
Building upon the legitimate variables identified above, we now examine the concept of conditional estimation parity, which pertains specifically to disparity in estimation error between groups distinguished by a protected attribute, conditional on legitimate features. Reducing the estimation error articulated in Eq. 4 thus mitigates the risk of conditional estimation disparity. However, directly assessing conditional estimation parity poses challenges due to the inherent difficulties in measuring estimation error.
In our case study, the Ombudsman presented evidence of the effects of the protected characteristics $x_{p}$ on the true prediction. For example, evidence showed that even when conditioned on legitimate variables, the model continued to rank Finnish-speaking and Swedish-speaking residents differently. This observation demonstrates a violation of conditional statistical parity (Eq. 3), as the predictions differ for groups distinguished by protected attributes after conditioning on legitimate variables. Additionally, because the Tribunal concluded there should be no legitimate differences in creditworthiness between these protected groups, the observed difference also indicates a violation of conditional estimation parity (Eq. 4). The difference arises specifically from the estimation error – the discrepancy between the true underlying parity (as established by the Tribunal) and the predictions generated by the model. The relevant legal evaluation is then whether it is a valid true difference or, as in this case, it is based on a protected characteristic $x_{p}$. Therefore, in this scenario, violations of conditional statistical parity and conditional estimation parity are closely related – the presence of estimation error directly leads to differential outcomes in model predictions. This highlights the importance of understanding the origins of prediction disparities through the lens of estimation errors, which can guide legal and practical evaluation of fairness and responsibility in predictive models.
Judicial legal reasoning is complex and requires engaging with this type of reasoning through statistical or theoretical means. Unlike other work that proposes quantitive thresholds or metrics for legality, we focus on formalising modelling choices within UK anti-discrimination law. This approach allows us to outline a flexible, decision-theoretic framework, aligning modeller and decision-maker choices as optimisation under legal constraints. Such a structured yet legally-informed framework aligns more closely with judicial reasoning, representing a fundamental and novel contribution to the literature on algorithmic discrimination.
@@ 8. Conclusion
This paper contributes to the understanding of unlawful discrimination in SML under UK law and related jurisdictions such as the EU and across the Commonwealth. Unlike the statistical focus in much of the algorithmic fairness literature, we present a legally grounded, decision-theoretic framework centred on the true data-generating process and its connection to legal causation.
By emphasising the legitimacy of prediction targets and features, we demonstrate how legally informed concepts, like conditional estimation parity, offer a more cohesive means of assessing when and how models discriminate. By situating these considerations within a UK context that shares key principles with many other jurisdictions, our approach enables more robust global approach to legal doctrine.
Crucially, our work corrects several mischaracterisations around discrimination, including those stemming from literature that misrepresents US law, fostering assumptions in the field that are not legally-informed, and literature that accurately reflects US law but diverges materially from UK law.